PQ is a smooth inclined plane whose inclination q can be varied in such a way that Q remains fixed and P can move on a vertical line PR. A particle slides from rest from P. For different values of q, the time of descent t from P is noted. Minimum value of t is 

Grade:12th pass

2 Answers

Arun
3 years ago
Force along the incline :
 
mg \sin \theta
 
Assuming zero Friction :
 
accel. = g \sin \theta
 
By Laws of motion, and that body starts from rest :
 
 {t}^{2} = \frac{2s}{g \sin \theta } 
 
And there you get the time taken by the body ^^
Vikas TU
3 years ago
Dear student 
The above explanation is not properly explained 
Please ignore this 
We are sorry for inconvenience.
Good Luck 
Cheers